# A New Benchmark and Model for Challenging Image Manipulation Detection

### v1

@@ Abstract
The ability to detect manipulation in multimedia data is vital in digital forensics. Existing Image Manipulation Detection (IMD) methods are mainly based on detecting anomalous features arisen from image editing or double compression artifacts. All existing IMD techniques encounter challenges when it comes to detecting small tampered regions from a large image. Moreover, compression-based IMD approaches face difficulties in cases of double compression of identical quality factors. To investigate the State-of-The-Art (SoTA) IMD methods in those challenging conditions, we introduce a new Challenging Image Manipulation Detection (CIMD) benchmark dataset, which consists of two subsets, for evaluating editing-based and compression-based IMD methods, respectively. The dataset images were manually taken and tampered with high-quality annotations. In addition, we propose a new two-branch network model based on HRNet that can better detect both the image-editing and compression artifacts in those challenging conditions. Extensive experiments on the CIMD benchmark show that our model significantly outperforms SoTA IMD methods on CIMD.
00footnotetext: *These authors contributed equally to this work.
@@ Introduction
Figure 1: Sample image of our dataset and comparison of image manipulation detection results with recent mainstream methods. The first three rows show manipulation of region copy-move, splicing and removal, respectively. The last row shows double-compressed splicing with the same Quality Factor (QF). Our method achieves the new state-of-the-art in detecting challenging manipulation cases.
Figure 2: Overview of the proposed two-branch architecture. RGB stream can detect anomalous features, while frequency stream is able to learn compression artifacts by feeding the image to the proposed compression artifacts learning model, as depicted in Fig. 5. The ASPP in Fig. 6(a) is appended to each of the outputs, and channel attention and spatial attention in Fig. 6(b)(c) interactively perform between each scale output to improve the detection performance under small manipulation.
There are three general types of image manipulation operations: (1) region splicing, where the content from one image is copied and pasted onto an image, (2) region copy-move, where an image regions is moved to another location within the same image, and (3) region removal, where parts of the image are erased and new contents are synthesized. To accurately detect these manipulations, some methods rely on detecting anomalous image region or texture features, while others identify double compression artifacts. While the State-of-the-Art (SoTA) IMD methods perform well on mainstream public IMD datasets, they still face two challenges as shown in Fig. 1. First, existing IMD methods have general difficulties in detecting relatively small tampered regions, due to the data-driven design under limited visual information. Secondly, approaches detecting double compression inconsistencies with two different quantization matrices fall apart when the compression Quality Factor (QF) remains the same. This is because the use of identical Q-matrix can significantly suppress double compression artifacts. As shown in Fig. 3, methods in this category detect tampered regions by identifying missing histogram values arisen from the two compression processes. When the same QF is used, the histogram undergoes very small changes, making it hard to detect double compression. In summary, as the image tampering techniques improve increasingly fast, forensic problems are typically ill-defined, and IMD methods in general fall behind in research for challenging cases.
@@
Figure 3: DCT coefficient histograms from the (0,1) position generated from a raw image under different compression processes.
@@
  • We present a novel two-branch architecture incorporating RGB and frequency features for challenging image manipulation detection. To our knowledge, our model is the first approach to focus on detecting small tampered regions.
@@
  • We introduce a new high-quality CIMD benchmark for evaluating the performance of SoTA IMD methods in challenging manipulations. We will make CIMD public upon paper acceptance.
@@
There are several datasets publicly available that are dedicated to image manipulation detection task. For example, the Columbia Dataset (Ng, Hsu, and Chang 2009) contains uncompressed 363 splicing images of a low average resolution ($938\times 720$). CASIA V1.0 and V2.0 (Dong, Wang, and Tan 2013a) were introduced for splicing-only manipulation detection with no ground truth mask. Numerous datasets have been introduced for copy-move tampering detection. For instance, the MICC (Amerini et al. 2011) features images mainly sourced from Columbia photographic image repository. Coverage (Wen et al. 2016) is another copy-move only dataset includes 100 original-forged pairs with similar-but-genuine objects. The NIST (Guan et al. 2019) has presented benchmark manipulation datasets with multiple versions. Some large benchmark datasets, such as (Mahfoudi et al. 2019) and (Novozamsky, Mahdian, and Saic 2020), apply non-realistic questionable automatically forgeries methods (Daisy et al. 2014) to generate forgery images. In addition, to detect compression artifacts, (Kwon et al. 2022) created five custom datasets that are double compressed using different unreported QFs.
Most existing datasets in image manipulation detection only focus on a specific type of manipulation or exhibit a significant imbalance in the distribution of tampered types. This results in unreliable measurement of a model’s overall detection capability. Furthermore, few datasets focus on challenging tampering detection. To address these limitations, we provide a novel dataset comprise two subsets: (1) Images have small regions of manipulation, with each tampering type containing an equal number of instances, and (2) Images with spliced double-compression using identical QFs. Details on our datasets and comparative analyses are in the supplementary material.
@@
In this work, we aim to build a comprehensive validation dataset (CIMD) dedicated to small region forgery (less than 1.5% on average) in both compressed and uncompressed scenarios. Our dataset are superior in dataset size, image quality, image diversity, and forgery strategy. Two separate subsets have been introduced to evaluate image editing-based and compression-based methods, respectively.
@@
Processing and Tampering. We used Photoshop 2023 (PS) to process and create tampering photos due to its popularity in other datasets mentioned in the related work section and its popularity in general public. More processing and tampering details are described in supplementary materials.
@@
The CIMD-R benchmark aims to provide a comprehensive evaluation of the image-editing-based models’ performance in detecting small tampered copy-move, object-removal, and splicing forgeries on uncompressed images. The use of uncompressed images eliminates undesired compression artifacts on forgery region that can be otherwise sensed by neural networks, enabling a more true performance evaluation on out-of-detection. CIMD-R comprises 600 TIFF images, with a resolution of 2048 $\times$ 1365. The ground-truth masks are also provided. In addition, CIMD-R adopts a future-oriented approach by providing 16-bit image pairs that offer up to $2^{48}$ (in trillions) colors. For copy-move manipulation, a part of an image is copied and pasted within the same image, followed by five post-processing methods that are scaling, rotation, level/curve increasing, illumination changing, and color redistribution. In the case of removal manipulation, forged images are synthesized by removing the selected region from the image (via Content-Aware Fill in PS). Content-Aware Fill is widely used in several datasets (Park et al. 2018b; Dong, Wang, and Tan 2013b) and represents the PS’s best guess to inpaint the object according to the surrounding region. For splicing forgery, regions from one image are copied and pasted into another source. Then, the same post-processing methods mentioned in copy-move are applied to make the forged region harmonious with its surroundings.
@@
The CIMD-C benchmark is designed to evaluate compressed-based model’s capability to detect double JEPG compression artifacts when the primary and secondary compression has the same QFs. The dataset comprises 200 JPEG images with a resolution of 2048 $\times$ 1365, wherein the QF is uniformly distributed as $10050\leq QF<$ 100. Forgery images are generated akin to SIMD-R’s splicing samples, with the distinction that the forged image is saved using the JPEG compression algorithm, employing the same QF as the original image. The original images were produced from RAW files ensuring that the original images are compressed for the first time, enhancing the dataset’s credibility. In the forgery images, the background is double-compressed, while the tampered regions are single-compressed. Furthermore, the dataset also comprises binary masks and QF values utilized for compression, thereby augmenting its utility for further investigations into the effects of different QFs.
@@
The two-branch architecture we propose enables the detection of both anomalous features and compression artifacts. Furthermore, our model is effective for detecting small manipulation regions and identifying double compression traces using the same quantization matrix (Q-matrix). To achieve our research objectives, we adopted HR-Net (Wang et al. 2020) as the backbone of our model, based on its ability to offer three-fold benefits. Firstly, the absence of pooling layers in HR-Net ensures that the features maintain high resolutions throughout the entire process. Secondly, the model processes features from different scales in parallel with effective information exchange, which is essential for capturing information of varying scales. Finally, the input size of HR-Net is ideally suited for DCT features. Since after processing by dilated convolution with a rate of 8, the size of the DCT feature is reduced to 1/8 of the input size, which is equivalent to the second stage resolution of HR-Net. We apply the full HR-Net for RGB streams, whereas for the frequency stream, we use the three-resolution variant HR-Net replacing the first stage with the proposed compression artifacts learning model shown in Fig. 5.
@@
The proposed network architecture comprises two branches, one for detecting anomalous features and the other for identifying compression artifacts, as in Fig. 2. For the RGB stream, the input image is fed to a full HR-Net, which learns the image editing traces from the visual content. In the frequency stream, the image is first input to the proposed compression artifact learning model shown in Fig. 5 to extract various DCT features. Subsequently, the DCT features are fed to a variant of the HR-Net, which operates at three different resolutions (1/8, 1/16, and 1/32).
To precisely pinpoint small tampering regions, we carefully designed our model using both Atrous Spatial Pyramid Pooling (ASPP) shown in Fig. 6(a) and Attention Mechanism. The ASPP captures long-range distance information via various receptive fields and handles scale variations. It consists of three dilated convolutional layers with different rates and a Global Average Pooling (GAP). The resulting features are concatenated and passed to a $1\times 1$ convolution.
@@
$R=\frac{1}{k}\sum_{i=1}^{k}(Q_{i}-Q_{i-1}).\vspace{-2mm}$ (5)
@@
Method Pixel-level F1 Image Level Best Fixed AUC Acc RRU-Net (Bi et al. 2019) 0.126 0.103 0.500 0.500 CR-CNN (Yang et al. 2020) 0.126 0.088 0.513 0.502 MantraNet (Wu et al. 2019) 0.051 0.018 0.500 0.500 SPAN (Hu et al. 2020) 0.160 0.045 0.510 0.498 PSCC-Net (Liu et al. 2022) 0.208 0.118 0.514 0.505 CAT-Net (Kwon et al. 2022) 0.301 0.194 0.589 0.537 MVSS-Net (Chen et al. 2021) 0.234 0.153 0.568 0.515 IF-OSN (Wu et al. 2022) 0.184 0.103 0.516 0.522 Ours 0.444 0.335 0.677 0.545
@@
$f(Q_{0}(i,j))=\left\{\begin{array}[]{ll}1,&\text{if}\left|clip(Q_{0}(i,j))\right|=t,t\in\left[0,T\right],\\ 0,&\text{otherwise}.\end{array}\right.\vspace{-0mm}$
The function $clip(\cdot)$ is utilized to extract the histogram feature within $[-T,T]$, which is essential for GPU memory constraints. We set $T$ as 20 from our experiments. Additionally, we apply the absolute operation as DCT histogram exhibits symmetry.
@@
Method Pixel-level F1 Image-level Optimal Fixed AUC Acc DJPEG (Park et al. 2018a) 0.026 0.022 0.500 0.500 Comprint (Mareen et al. 2022) 0.030 0.010 0.467 0.500 CAT-Net (Kwon et al. 2022) 0.395 0.259 0.534 0.490 Ours 0.542 0.442 0.727 0.525
In order to capture the inter-block relationship of each frequency, the binary volume features are input to an atrous convolution with a dilated rate of 8. The dilated convolution (Yu and Koltun 2015) was originally designed to improve receptive fields without downsampling. The features are then input to $1\times 1$ convolution for channel-reducing purposes. The compression artifact learning method involves two element-wise multiplication operations. The first multiplication is performed between the histogram features and the Q-matrix, which is utilized to simulate the JPEG de-quantization procedure. The second multiplication is used to guide the histogram feature to focus more on unstable coefficients, which is a critical step for detecting double-compressed images using the same QF.
@@
Datasets. The training datasets used in this study were mostly adopted from (Kwon et al. 2022), which comprise CASIA v2, Fantastic Reality, IMD2020, and the datasets specifically designed for detecting compression artifacts using different QFs. The testing phase entailed the utilization of CIMD-R and CIMD-C to evaluate the efficacy of image-editing-based and compression-based methods, respectively. Further details pertaining to the datasets employed and a comparative analysis of evaluation metrics with select publicly accessible datasets are provided in the supplementary materials.
Evaluation metrics. Following most previous work, we evaluated the localization results using pixel-level F1 score with both the optimal and fixed 0.5 thresholds. For image-level detection, we employed AUC and image-level accuracy. We also set 0.5 as the threshold for image-level accuracy. Only tampered images are used for the manipulation localization evaluation.
Implementation details. Our model was implemented using PyTorch (Paszke et al. 2019) and trained on 8 RTX 2080 GPUs, with batch size 4. We set the initial learning rate as 0.001 with exponential decay. To mitigate the impact of imbalanced datasets for model training, we randomly selected an equal number of samples from each dataset for each epoch. The training process consists of 250 epochs. The proposed model is designed to accept various image formats, including both JPEG and non-JPEG formats. For non-JPEG images, the model employs an all-1 Q-matrix to compress the samples, which is equivalent to using a QF of 100 for lossless compression. The backbone of RGB stream is pre-trained using ImageNet (Krizhevsky, Sutskever, and Hinton 2017), while the frequency stream is pre-trained using double-compressed images introduced by (Park et al. 2018a). To enhance the model’s sensitivity in detecting small tampering regions, the training objective is designed to minimize the pixel-level binary cross-entropy loss.
@@
To guarantee a fair comparison and evaluate the previous models using newly introduced CIMD, we select the state-of-the-art approaches using these two standards: (1) pre-trained model is publicly available, and (2) the evaluation datasets we used are not in their training sets. Following these criteria, we select RRU-Net, MantraNet , CR-CNN, SPAN, PSCC-Net, MVSS-Net, IF-OSN, CAT-Net, DJPEG and Comprint. Among them, DJPEG and Comprint are designed for compression artifacts detection, while CAT-Net can detect anomalous feature and compression artifacts jointly. All the aforementioned studies are appropriately referenced in the related work section. We use CIMD-R to evaluate the performance of the image-editing-based method, while CIMD-C is utilized for compression-based approaches.
Evaluation using CIMD-R subset. Table 1 reports the results of image-editing-based methods using CIMD-R, in which all image samples are uncompressed. Pixel-level F1 score is calculated using the best F1 threshold for each image and using fixed F1 threshold of 0.5. Best scores are highlighted in bold. Our approach achieves the best performance in both image-level and pixel-level detection tasks. Notably, our method outperforms existing SoTA methods in both image-level and pixel-level evaluation, which demonstrates its superiority for detecting small tampering regions.
@@
Ablation study. We provide a simple ablation study shown in Table 3. Observe that our RGB stream is effective in both compressed and uncompressed data. Notably, the frequency stream fails to produce satisfactory results in CIMD-R due to the absence of compression artifacts. However, when the two branches work collaboratively, the model’s performance improves in both localization and detection evaluation. Additional ablation studies and experimental results are provided in the supplementary material.
Table 3: Ablation study of two streams to work collaboratively and/or separately. Pixel-level F1 using 0.5 and image-level AUC are reported.

### v2

@@ Abstract
The ability to detect manipulation in multimedia data is vital in digital forensics. Existing Image Manipulation Detection (IMD) methods are mainly based on detecting anomalous features arisen from image editing or double compression artifacts. All existing IMD techniques encounter challenges when it comes to detecting small tampered regions from a large image. Moreover, compression-based IMD approaches face difficulties in cases of double compression of identical quality factors. To investigate the State-of-The-Art (SoTA) IMD methods in those challenging conditions, we introduce a new Challenging Image Manipulation Detection (CIMD) benchmark dataset, which consists of two subsets, for evaluating editing-based and compression-based IMD methods, respectively. The dataset images were manually taken and tampered with high-quality annotations. In addition, we propose a new two-branch network model based on HRNet that can better detect both the image-editing and compression artifacts in those challenging conditions. Extensive experiments on the CIMD benchmark show that our model significantly outperforms SoTA IMD methods on CIMD. The dataset is available at: https://github.com/ZhenfeiZ/CIMD.
@@ Introduction
Figure 1: Sample images of our dataset and comparison of image manipulation detection results with recent mainstream methods. The first three rows show manipulation of region copy-move, splicing and removal, respectively. The last row shows double-compressed splicing with the same Quality Factor (QF). Our method achieves the new state-of-the-art in detecting challenging manipulation cases.
Figure 2: Overview of the proposed two-branch architecture. RGB stream can detect anomalous features, while frequency stream is able to learn compression artifacts by feeding the image to the compression artifacts learning model, as depicted in Fig. 5. The ASPP in Fig. 6(a) is appended to each of the outputs, and channel attention and spatial attention in Fig. 6(b)(c) interactively perform between each scale output to improve the detection performance under small manipulation.
There are three general types of image manipulation operations: (1) region splicing, where the content from one image is copied and pasted onto another image, (2) region copy-move, where an image regions is moved to another location within the same image, and (3) region removal, where parts of the image are erased and new contents are synthesized. To accurately detect these manipulations, some methods rely on detecting anomalous image region or texture features, while others identify double compression artifacts. While the State-of-the-Art (SoTA) IMD methods perform well on mainstream public IMD datasets, they still face two challenges as shown in Fig. 1. First, existing IMD methods have general difficulties in detecting relatively small tampered regions, due to the data-driven design under limited visual information. Secondly, approaches detecting double compression inconsistencies with two different quantization matrices fall apart when the compression Quality Factor (QF) remains the same. This is because the use of identical Q-matrix can significantly suppress double compression artifacts. As shown in Fig. 3, methods in this category detect tampered regions by identifying missing histogram values arisen from the two compression processes. When the same QF is used, the histogram undergoes very small changes, making it hard to detect double compression. In summary, as the image tampering techniques improve increasingly fast, forensic problems are typically ill-defined, and IMD methods in general fall behind in research for challenging cases.
@@
Figure 3: DCT coefficient histograms from the (0,1) position generated from a raw image under different compression processes. The range of X-axis is [-20, 20].
@@
  • We present a two-branch architecture incorporating RGB and frequency features for challenging image manipulation detection. To our knowledge, our model is the first approach to focus on detecting small tampered regions.
@@
  • We introduce a new high-quality CIMD benchmark for evaluating the performance of SoTA IMD methods in challenging manipulations.
@@
There are several datasets publicly available that are dedicated to image manipulation detection task. For example, the Columbia Dataset (Ng, Hsu, and Chang 2009) contains uncompressed 363 splicing images of a low average resolution ($938\times 720$). CASIA V1.0 and V2.0 (Dong, Wang, and Tan 2013a) were introduced for splicing and copy-move manipulation detection with no ground truth mask. Numerous datasets have been introduced only for copy-move tampering detection. For instance, the MICC (Amerini et al. 2011) features images mainly sourced from Columbia photographic image repository. Coverage (Wen et al. 2016) is another copy-move only dataset includes 100 original-forged pairs with similar-but-genuine objects. The NIST (Guan et al. 2019) has presented benchmark manipulation datasets with multiple versions. Some large benchmark datasets, such as (Mahfoudi et al. 2019) and (Novozamsky, Mahdian, and Saic 2020), apply non-realistic questionable automatically forgeries methods (Daisy et al. 2014) to generate forgery images. In addition, to detect compression artifacts, (Kwon et al. 2022) created five custom datasets that are double compressed using different unreported QFs.
Most existing datasets in image manipulation detection only focus on a specific type of manipulation or exhibit a significant imbalance in the distribution of tampered types. This results in unreliable measurement of a model’s overall detection capability. Furthermore, few datasets focus on challenging tampering detection. To address these limitations, we provide a novel dataset comprise two subsets: (1) Images with small manipulation regions, where each tampering type contains an equal number of instances, and (2) Images with spliced double-compression using identical QFs.
@@
In this work, we aim to build a comprehensive validation dataset (CIMD) dedicated to small region forgery (less than 1.5% on average) in both compressed and uncompressed scenarios. Our dataset are superior in image quality, image diversity, and forgery strategy. Two separate subsets have been introduced to evaluate image editing-based and compression-based methods, respectively.
@@
Processing and Tampering. We used Photoshop 2023 (PS) to process and create tampering photos due to its popularity in other datasets mentioned in the related work section and its popularity in general public.
@@
The CIMD-R benchmark provides a comprehensive evaluation of the image-editing-based models’ performance in detecting small tampered copy-move, object-removal, and splicing forgeries on uncompressed images. The use of uncompressed images eliminates undesired compression artifacts on forgery region that can be otherwise sensed by neural networks, enabling a more true performance evaluation on out-of-detection. CIMD-R comprises 600 TIFF images, with a resolution of 2048 $\times$ 1365. Ground-truth masks are also provided. In addition, CIMD-R adopts a future-oriented approach by providing 16-bit image pairs that offer up to $2^{48}$ (trillions of) colors. For copy-move manipulation, a part of an image is copied and pasted within the same image, followed by five post-processing methods: scaling, rotation, level/curve increasing, illumination changing, and color redistribution. In the case of removal manipulation, forged images are synthesized by removing the selected region from the image (via Content-Aware Fill in PS). Content-Aware Fill is widely used in several datasets (Park et al. 2018b; Dong, Wang, and Tan 2013b) and represents the PS’s best guess to inpaint the object according to the surrounding region. For splicing forgery, regions from one image are copied and pasted into another source. Then, the same post-processing methods mentioned in copy-move are applied to make the forged region harmonious with its surroundings.
@@
The CIMD-C benchmark is designed to evaluate the capability of compressed-based models in detecting double JEPG compression artifacts, where the primary and secondary compression has the same QFs. The dataset comprises 200 JPEG images with a resolution of 2048 $\times$ 1365, wherein the QF is uniformly distributed as $10050\leq QF<$ 100. Forgery images are generated akin to CIMD-R’s splicing samples, with the distinction that the forged image is saved using the JPEG compression algorithm, employing the same QF as the original image. The original images were produced from RAW files ensuring that the original images are compressed for the first time, enhancing the dataset’s credibility. In the forgery images, the background is double-compressed, while the tampered regions are single-compressed. Furthermore, the dataset also comprises binary masks and QF values utilized for compression, thereby augmenting its utility for further investigations into the effects of different QFs.
@@
The two-branch architecture we propose enables the detection of both anomalous features and compression artifacts inspired by (Kwon et al. 2022). Furthermore, our model is effective for detecting small manipulation regions and identifying double compression traces that apply the same quantization matrix (Q-matrix). To achieve our research objectives, we adopted HR-Net (Wang et al. 2020) as the backbone of our model, based on its ability to offer three-fold benefits. Firstly, the absence of pooling layers in HR-Net ensures that the features maintain high resolutions throughout the entire process. Secondly, the model processes features from different scales in parallel with effective information exchange, which is essential for capturing information of varying scales. Finally, the input size of HR-Net is ideally suited for DCT features. Since after processing by dilated convolution with a rate of 8, the size of the DCT feature is reduced to 1/8 of the input size, which is equivalent to the second stage resolution of HR-Net.
@@
The network architecture comprises two branches, one for detecting anomalous features and the other for identifying compression artifacts, as in Fig. 2. For the RGB stream, the input image is fed to a full HR-Net, which learns the image editing traces from the visual content. In the frequency stream, the image is first input to the proposed compression artifact learning model shown in Fig. 5 to extract various DCT features. Subsequently, the DCT features are fed to a variant of the HR-Net, which operates at three different resolutions (1/8, 1/16, and 1/32).
To precisely pinpoint small tampering regions, we carefully designed our model using both Atrous Spatial Pyramid Pooling (ASPP) shown in Fig. 6(a) and Attention Mechanism shown in Fig. 6(b)(c). The ASPP captures long-range distance information via various receptive fields and handles scale variations. It consists of three dilated convolutional layers with different rates and a Global Average Pooling (GAP). The resulting features are concatenated and passed to a $1\times 1$ convolution.
@@
$R=\frac{1}{k}\sum_{i=1}^{k}(Q_{i}-Q_{i-1}).$ (5)
Method Pixel-level F1 Image Level Best Fixed AUC Acc RRU-Net (Bi et al. 2019) 0.126 0.103 0.500 0.500 CR-CNN (Yang et al. 2020) 0.126 0.088 0.513 0.502 MantraNet (Wu et al. 2019) 0.051 0.018 0.500 0.500 SPAN (Hu et al. 2020) 0.160 0.045 0.510 0.498 HiFi_IFDL (Guo et al. 2023) 0.145 0.115 0.502 0.502 PSCC-Net (Liu et al. 2022) 0.208 0.118 0.514 0.505 CAT-Net (Kwon et al. 2022) 0.301 0.194 0.589 0.537 MVSS-Net (Chen et al. 2021) 0.234 0.153 0.568 0.515 IF-OSN (Wu et al. 2022) 0.184 0.103 0.516 0.522 Ours 0.444 0.335 0.677 0.545
@@
$f(Q_{0}^{t}(i,j))=\left\{\begin{array}[]{ll}1,&\text{if}\left|clip(Q_{0}(i,j))\right|=t,t\in\left[0,T\right],\\ 0,&\text{otherwise}.\end{array}\right.$
The function $clip(\cdot)$ is utilized to extract the histogram feature within $[-T,T]$, which is essential for GPU memory constraints. We set $T$ as 20 from the experiments. Additionally, we apply the absolute operation as DCT histogram exhibits symmetry.
Method Pixel-level F1 Image-level Best Fixed AUC Acc DJPEG (Park et al. 2018a) 0.026 0.022 0.500 0.500 Comprint (Mareen et al. 2022) 0.030 0.010 0.467 0.500 CAT-Net (Kwon et al. 2022) 0.395 0.259 0.534 0.490 Ours 0.542 0.442 0.727 0.525
@@
The compression artifact learning method involves two element-wise multiplication operations. The first multiplication is performed between the histogram features and the Q-matrix, which is utilized to simulate the JPEG de-quantization procedure. The second multiplication is used to guide the histogram feature to focus more on unstable coefficients, which is a critical step for detecting double-compressed images using the same QF.
@@
Datasets. The training datasets used in this study were adopted from (Kwon et al. 2022). The testing phase entailed the utilization of CIMD-R and CIMD-C to evaluate the efficacy of image-editing-based and compression-based methods, respectively.
Evaluation metrics. Following most previous work, we evaluated the localization results using pixel-level F1 score with both the optimal and fixed 0.5 thresholds. For image-level detection, we employed AUC and image-level accuracy. We set 0.5 as the threshold for image-level accuracy. Only tampered images are used for the manipulation localization evaluation.
Implementation details. Our model was implemented using PyTorch (Paszke et al. 2019) and trained on 8 RTX 2080 GPUs, with batch size 4. We set the initial learning rate as 0.001 with exponential decay. The training process consists of 250 epochs. The proposed model is designed to accept various image formats, including both JPEG and non-JPEG formats. The training objective is designed to minimize the pixel-level binary cross-entropy.
@@
To guarantee a fair comparison and evaluate the previous models using newly introduced CIMD, we select the state-of-the-art approaches using these two standards: (1) pre-trained model is publicly available, and (2) the evaluation datasets we used are not in their training sets. Following these criteria, we select RRU-Net, MantraNet, HiFi_IFDL, CR-CNN, SPAN, PSCC-Net, MVSS-Net, IF-OSN, CAT-Net, DJPEG and Comprint. All the work we compared are appropriately referenced in the related work section. We use CIMD-R to evaluate the performance of the image-editing-based method, while CIMD-C is utilized for compression-based approaches.
Evaluation using CIMD-R subset. Table 1 reports the results of image-editing-based methods using CIMD-R, in which all image samples are uncompressed. Two Pixel-level F1 scores are calculated using the best F1 threshold for each image and using fixed F1 threshold of 0.5, respectively. Best scores are highlighted in bold. Our method outperforms existing SoTA methods in both image-level and pixel-level evaluation, which demonstrates its superiority for detecting small tampering regions.
@@
Ablation study. We provide a simple ablation study shown in Table 3. Observe that our RGB stream is effective in both compressed and uncompressed data. Notably, the frequency stream fails to produce satisfactory results in CIMD-R due to the absence of compression artifacts. However, when the two branches work collaboratively, the model’s performance improves in both localization and detection evaluation.
@@
Table 3: Ablation study of two streams to work collaboratively and/or separately.